# Can LLMs Generate Diverse Molecules? Towards Alignment with Structural Diversity

### v1

@@ 4.4 Ablation studies
Div-SFT+RL 21.98 16.98 14.35 0.45 0.74
Large number of samples vs. performance. We also analyze how well our method discovers diverse molecules with respect to the number of generations. In this experiment, we extend beyond the settings in Section 4.1. We consider our method to generate $85$, $120$, and $155$ molecules with a single NVIDIA A100 SXM4 40GB GPU. Next, we consider beam search with beam sizes of $300$, $400$, and $500$, respectively. The beam search is implemented with four NVIDIA A100 SXM4 40GB GPUs due to the memory limitation of a single GPU. Note that this experiment uses $250$ molecular descriptions in the ChEBI-20 test set. We present the results in Table 3. One can see that our method (1) discovers more diverse molecules with respect to the total number of generations and (2) exhibits further performance improvements as the number of generations increases.
Time costs vs. performance. In addition, we also analyze how well our method discovers diverse molecules with respect to the time costs. In Table 3, we present the time costs for each method. One can see that our method discovers more diverse molecules in a practical time compared to the decoding schemes for diverse sequence generation.
SFT with hard filtering vs. RL. To train LLMs to generate diverse molecules, one may also consider supervised fine-tuning on hard-filtered datasets, e.g., using a set of molecules filtered by similarity, as an alternative to reinforcement learning. In this experiment, we additionally perform supervised fine-tuning with distinct molecules, where each pair of molecules yields a Tanimoto similarity below $0.65$. The results are presented in Table 4. One can see that while it yields relatively high NCircles with respect to the number of accepted unique molecules, the overall performance is worse compared to the performance of reinforcement learning.
5 Conclusion
In this paper, we identify the limitations of large language models (LLMs) for generating diverse molecules. Then, we present new supervised fine-tuning and reinforcement learning methods to adapt existing molecular generative LLMs to generate a diverse set of molecules. Experiments show that our approach enables LLMs to better discover diverse molecules compared to the existing approaches. An interesting avenue for future work is to reduce the space and time complexity in generating the sequence of molecules, for example, based on set encoding techniques (Zaheer et al., 2017).
Appendix A Detailed implementations and training
A.1 Supervised fine-tuning
Dataset Collection. To fine-tune BioT5+ and MolT5 (Sections 4.1 and 4.2), we collect $T=100$ molecules using beam search for each training molecular description in the ChEBI-20 training dataset. Note that this dataset has been considered in the original BioT5+ and MolT5.
[…]
Supervised learning. We consider four NVIDIA A100 GPUs for supervised fine-tuning.
  • For the supervised fine-tuning of BioT5+ and MolT5 (Sections 4.1 and 4.2), we consider $80$ epochs, $8$-batch size, $5\textrm{e}-4$ learning rate, $0.05$ warm-up ratio, and apply a cosine learning scheduler. The maximum sequence length in supervised training is limited to $2560$ due to memory limitations.
@@ A.1 Supervised fine-tuning
A.2 Reinforcement learning
Reward Design. In experiments with BioT5+ and MolT5 on the ChEBI-20 dataset (Sections 4.1 and 4.2), we define the description-matching reward using the BLEU score as follows:
$r_{\text{match}}(m_{k},p_{\text{desc}})=\text{BLEU}(m_{p_{\text{mol}}},m_{k})^{\alpha},$ (2)
where $m_{p_{\text{mol}}}$ is a ground-truth molecule satisfying the given description. The ChEBI-20 dataset involves a set of pairs $(p_{\text{desc}},m_{p_{\text{mol}}})$. Note that $\alpha$ is a hyper-parameter.
For experiments with DrugAssist (Section 4.3), the property-matching reward yields $1$ if the molecule satisfies the quantitative properties described in $p_{\text{desc}}$, e.g., HB donors, HB acceptors, and Bertz complexity, and 00 otherwise. The properties are evaluated with the RDKit package (Landrum, 2016).
Next, the diversity reward, $r_{\text{div}}$, is defined to consider molecular structural diversity as follows:
@@ A.2 Reinforcement learning
where $T(m_{i},m_{j})$ measures the Tanimoto similarity between $m_{i}$ and $m_{j}$ by assessing their true molecular structures, i.e., fingerprints. This metric yields a value between 00 and $1$ and can be obtained using the RDKit package. Note that $\beta$ is a hyper-parameter.
@@
  • For the reinforcement learning of BioT5+ and MolT5 (Sections 4.1 and 4.2), we consider $200$ PPO iterations, $8$ mini-batch size, $128$ batch size, and $5\textrm{e}-5$ learning rate. We also consider $0.01$ KL penalty. Note that $\alpha$ and $\beta$ in Equations 2 and 3 are $0.5$ and $2.0$, respectively. The reward signal is amplified by multiplying by a value of $8.0$. The maximum sequence length in reinforcement learning is limited to $2560$ due to memory limitations. Here, we also apply LoRA (Hu et al., 2022) where the rank and alpha are $16$ and $32$, respectively. We save the model every $25$ PPO iteration and select the model yielding the highest rewards for the training prompts.
  • For the reinforcement learning of DrugAssist (Section 4.3), we consider $200$ PPO iterations, $4$ mini-batch size, $64$ batch size, and $3\textrm{e}-6$ learning rate. We also consider $0.1$ KL penalty. Note that $\beta$ in Equation 3 is $2.0$. The reward signal is amplified by multiplying by a value of $4.0$. The maximum sequence length in reinforcement learning is limited to $1280$. We also apply LoRA (Hu et al., 2022) where the rank and alpha are $64$ and $128$, respectively. We save the model every $25$ PPO iteration and select the model yielding the highest rewards for the training prompts.
Appendix B Detailed experimental settings
Comparison with decoding schemes (Section 4.1). In this experiment, we first consider random sampling with different temperatures $\{0.7,1.0,1.5\}$. For the other decoding schemes, we consider conventional configurations: nucleus sampling with top-p $0.8$, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. We apply greedy decoding for our approach. The prompts for BioT5+ are described in Table 5. The prompts for MolT5 are considered as a molecular description without any additional comments.
@@ Appendix B Detailed experimental settings
Comparison with LLMs (Section 4.2). In this experiment, we consider generalist LLMs. For Mol-instructions, and LlasMol, we apply random sampling, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. The prompts are described in Table 6. For ChatGPT, we apply either $50$ random sampling with $p_{\text{desc}}$ or greedy decoding with $p_{\text{desc+div}}$, where both prompts are described in Table 7. The prompt design is based on (Li et al., 2024). For o1, we use the prompt $p_{\text{desc+div}}$ described in Table 7. We report the test score using the decoding scheme or the prompt that yields the highest NCirclesh=0.65 for $500$ evaluation molecular description.
Table 6: Prompts for generalist LLMs (Fang et al., 2024; Taylor et al., 2022).
@@
Mol-instructions “Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: Create a molecule with the structure as the one describe. ### Input: <molecular description> ### Response: ”
@@
Table 7: Prompts for ChatGPT and o1 (Li et al., 2024).
Input Contents
System $p_{\text{desc}}$ “You are now working as an excellent expert in chemistry and drug discovery. Given the caption of a molecule, your job is to predict the SMILES representation of the molecule. The molecule caption is a sentence that describes the molecule, which mainly describes the molecule’s structures, properties, and production. You can generate the molecule SMILES representation from the caption. Example: Instruction: Given the caption of a molecule, predict the SMILES representation of the molecule. Input: <molecular description>. Your output should be: {"molecule": "<SMILES>"} Your response should only be in the JSON format above; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.”
User $p_{\text{desc}}$ “Input: <molecular description>”
System $p_{\text{desc+div}}$ “You are now working as an excellent expert in chemistry and drug discovery. Given the caption of a molecule, your job is to generate diverse molecules. The molecule caption is a sentence that describes the molecule, which mainly describes the molecule’s structures, properties, and production. You can generate the molecule SMILES representation from the caption. Example: Instruction: Given the caption of a molecule, generate fifty structurally diverse molecules with SMILES representations. Input: <molecular description>. Your output should be: {"molecule1": "<SMILES>"} {"molecule2": "<SMILES>"} {"molecule3": "<SMILES>"} {"molecule50": "<SMILES>"} Your response should only be in the JSON format above; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.”
User $p_{\text{desc+div}}$ “Input: <molecular description>”
Fine-tuning generalist LLMs. We synthesize $600$ pairs of training prompts and corresponding sets of molecules. The prompts specify hydrogen bond donors and acceptors ranging from one to four, and a Bertz complexity ranging from 00 to $300$. The sets of molecules are collected by applying beam search on DrugAssist and then perturbed by shuffling their order. We use the prompts described in Table 8. The system prompt follows the default settings from (Ye et al., 2023). For evaluation, we use four prompts specifying three hydrogen bond donors, three hydrogen bond acceptors, a Bertz complexity between $100$ and $200$, and a QED value between $0.4$ and $0.6$. Additionally, as shown in Figure 3, we try to generate diverse molecules by designing prompts without fine-tuning (Table 9). However, these prompts show lower performance compared to applying beam search.
@@
“Can you generate fifty molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate fifty diverse molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate fifty structurally diverse molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
@@
“Can you generate fifty molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate fifty diverse molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate fifty structurally diverse molecules with <value> <property>? Print each of them in SMILES format.”
Appendix C Additional results
Table 10: Visualization of the generated molecules from beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.53$)
Table 11: Visualization of the generated molecules from diverse beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.67$)
Table 12: Visualization of the generated molecules from constrastive beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.58$)
Table 13: Visualization of the generated molecules from our method. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F.”
Example of outputs (IntDiv. is $0.77$)

### v2

@@ A.2 Molecular Similarity Measures
A.3 Molecular Diversity Metrics
In this section, we provide a detailed explanation of the diversity metrics for evaluating the given set of molecules. Specifically, we explain two diversity metrics: the number of circles (Xie et al., 2023) and the internal diversity (Polykovskiy et al., 2020).
The number of circles (NCircles.; Xie et al., 2023). To evaluate the diversity of a given set of molecules $\mathcal{M}$, this computes the size of the largest subset of molecules in which no two molecules are similar to each other. This metric is defined with a Tanimoto similarity $T(\cdot,\cdot)$ and a similarity threshold $h$ as follows:
${\text{NCircles}_{h}=\text{max}_{\mathcal{C}\subseteq\mathcal{M}}\;|\mathcal{C}|\quad\text{s.t. }T(x,y)<h,\forall x\neq y\in\mathcal{C},}$ (2)
where $\mathcal{C}$ is a subset of molecules. Every pair of molecules in $\mathcal{C}$ should have a similarity lower than $h$. The high NCircles value implies that the given set of molecules $\mathcal{M}$ is diverse and covers a wide range of molecular space (Xie et al., 2023). Recent work (Renz et al., 2024) haven shown that this metric is relatively exact to measure the molecular diversity, compared to the other metrics, e.g., internal diversity.
Internal diversity (IntDiv.; Polykovskiy et al., 2020). Given a set of molecules $\mathcal{M}$, this metric measures the average of pair-wise Tanimoto similarities:
@@ A.3 Molecular Diversity Metrics
where $m_{i}$ is $i$-th molecule in the given set of molecules $\mathcal{M}$.
Appendix B Detailed Implementations and Training
B.1 Supervised Fine-tuning
Dataset collection. To fine-tune BioT5+ (Section 4.1), we collect $T=100$ molecules using contrastive beam search for each training molecular description in ChEBI-20 training split. In the case of molecular descriptions from the L+M-24 training split, we use only the provided set of target molecules for each description, without additional data collection. To fine-tune DrugAssist (Section 4.3), we collect $T=300$ molecules using contrastive beam search for each training prompt. Note that the collected molecules were filtered to remove invalid string representations, duplicated molecules, and unaccepted molecules. The invalid string representations are evaluated with RDKit package (Landrum, 2016). Additionally, the collected molecules are concatenated into a single sequence $m_{1}||\cdots||m_{K}$.
Supervised learning. We consider four NVIDIA A100 SXM4 40GB GPUs for supervised fine-tuning.
  • For the supervised fine-tuning of BioT5+ (Section 4.1), we consider $80$ epochs, $8$-batch size, $5\textrm{e}-4$ learning rate, $0.05$ warm-up ratio, and apply a cosine learning scheduler. The maximum sequence length in supervised training is limited to $2560$ due to memory limitations.
@@ B.1 Supervised Fine-tuning
B.2 Reinforcement Learning
Reward design. In experiments on L+M-24 and ChEBI-20 datasets (Section 4.1), we define the description-matching reward using the Dice similarity as follows:
$r_{\text{match}}(m_{k},p_{\text{desc}})=\max_{m_{\text{target}\in\mathcal{M}_{\mathcal{D}}(p_{\text{desc}})}}D(m_{\text{target}},m_{k})^{\alpha},$ (4)
where $\mathcal{M}_{\mathcal{D}}(p_{\text{desc}})$ is a set of target molecules provided in the datasets for each molecular description $p_{\text{desc}}$. [^7] As described in Section 4.1, we assume that the molecule satisfies the description if it shares a certain degree of structures with one of the target molecules provided in the dataset. As described in Section A.2, $D(m_{i},m_{j})$ is Dice similarity captuing the degree of shared structural features between two molecules. Note that $\alpha$ is a hyper-parameter. In experiments with DrugAssist (Section 4.3), the description-matching reward yields $1$ if the molecule satisfies the quantitative properties described in $p_{\text{desc}}$ (evaluated using RDkit (Landrum, 2016)) and 00 otherwise.
Next, the diversity reward, $r_{\text{div}}$, is defined to consider molecular structural diversity between the new molecule $m_{k}$ and the previously generated molecules $\{m_{i}\}^{k-1}_{i=1}$ within a sequence:
@@ B.2 Reinforcement Learning
where $T(m_{i},m_{j})$ is Tanimoto similarity (Section A.2) between $m_{i}$ and $m_{j}$ and $\beta$ is a hyper-parameter.
@@
  • For the reinforcement learning of BioT5+ (Section 4.1), we consider $200$ PPO iterations, $8$ mini-batch size, $128$ batch size, and $5\textrm{e}-5$ learning rate. We also consider $0.01$ KL penalty. Note that $\alpha$ and $\beta$ in Equations 4 and 5 are $0.5$ and $1.0$, respectively. We increate $\beta$ to $2.0$ during training on ChEBI-20 dataset. The reward signal is amplified by multiplying by a value of $8.0$. The maximum sequence length in reinforcement learning is limited to $2560$ due to memory limitations. Here, we also apply LoRA (Hu et al., 2022) where the rank and alpha are $16$ and $32$, respectively.
  • For the reinforcement learning of DrugAssist (Section 4.3), we consider $200$ PPO iterations, $4$ mini-batch size, $64$ batch size, and $3\textrm{e}-6$ learning rate. We also consider $0.1$ KL penalty. Note that $\beta$ in Equation 5 is $2.0$. The reward signal is amplified by multiplying by a value of $8.0$. The maximum sequence length in reinforcement learning is limited to $1024$ due to memory limitations. We also apply LoRA (Hu et al., 2022) where the rank and alpha are $64$ and $128$, respectively.
Appendix C Dataset Details
Table 7: Data statistics. The number of molecular descriptions.
Dataset $\sharp$ of training $\sharp$ of test
L+M-24 160,492 21,839
ChEBI-20 26,408 3,301
In this section, we describe detailed data statistics of ChEBI-20 (Edwards et al., 2021) and L+M-24 datasets (Edwards et al., 2024). Both datasets serve to validate description-guided molecular generation capability. Note that both datasets consider a molecular description expressed in English and a molecule represented by SMILES (Weininger, 1988). In Table 7, we provide the number of molecular descriptions.
To be specific, L+M-24 dataset provides an average of $70$ and $17$ target molecules for each description in training and test splits, respectively. Note that the original L+M-24 dataset provides a single target molecule for each description. However, L+M-24 dataset involves identical molecular properties for different molecules. Moreover, some molecular descriptions, e.g., an inhibitor of both BCL2 and BTK proteins, include the molecular properties of other descriptions, e.g., an inhibitor of BCL2 protein. Thus, we associate multiple molecules with a single molecular description. Next, ChEBI-20 dataset provides a single example of target molecule for each description.
In addition, since the original BioT5+ is not pre-trained on L+M-24 dataset, we further train it on the original L+M-24 training split. Note that we use the first $1,000$ molecular descriptions in the test split of the L+M-24 dataset for evaluation, as evaluating the entire test split with $50$ generated molecules per description takes too much time (five to six days).
Appendix D Experiments Setup
Our experiments consider various molecular generative LLMs: MolT5 (Edwards et al., 2022), BioT5+ (Pei et al., 2024), Text+Chem T5 (Christofidellis et al., 2023), LlasMol (Yu et al., 2024), DrugAssist (Ye et al., 2023), and Meditron (Chen et al., 2023). [^8] MolT5, BioT5+, Text+Chem T5, LlasMol, and DrugAssist are licensed under the MIT License. [^9] Meditron is licensed under the Apache 2.0 License. They consider a molecular description expressed in English. They also consider a molecule represented as SMILES (Weininger, 1988), except for BioT5+, which uses SELFIES (Krenn et al., 2020). T5 models have 220M parameters, and others have 7B parameters. We use four NVIDIA A100 SXM4 40GB GPUs for the experiments. We consider a single run.
Comparison with decoding schemes. In this experiment, we first consider random sampling with different temperatures $\{0.7,1.0,1.5\}$. For the other decoding schemes, we consider conventional configurations: nucleus sampling with top-p $0.8$, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. We apply greedy decoding for our approach. The prompts for BioT5+ are described in Table 8.
@@ Appendix D Experiments Setup
Comparison with existing LLMs. For all existing LLMs, we apply random sampling, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. The prompts for existing LLMs, which are molecular generative LLMs based on a given molecular description, are described in Table 9.
Table 9: Prompts for various existing LLMs
@@
MolT5 “<molecular description>”
Meditron “Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: You are a researcher. You can come up molecule smile strings based on your existing knowledge. Molecule smile strings are given against the following input. You should be as detailed as possible. Input: <molecular description> In that caption, could you generate a molecule smile string? ### Response: ”
Text+Chem T5 “<molecular description>”
@@
Fine-tuning generalist LLMs. We synthesize $600$ pairs of training prompts and corresponding sets of molecules. The prompts specify hydrogen bond donors and acceptors ranging from one to four, and a Bertz complexity ranging from 00 to $300$. The sets of molecules are collected by applying beam search on DrugAssist and then perturbed by shuffling their order. We use the prompts described in Table 10. The system prompt follows the default settings from (Ye et al., 2023). For evaluation, we use four prompts specifying three hydrogen bond donors, three hydrogen bond acceptors, a Bertz complexity between $100$ and $200$, and a QED value between $0.4$ and $0.6$. Additionally, as shown in Figure 1, we try to generate diverse molecules by designing prompts (Table 11). However, these prompts show lower performance compared to applying beam search.
Appendix E Additional Results
In this section, we present qualitative results by evaluating whether the generated molecules satisfy the given molecular description using an external tool. Here, we consider the generation of a set of molecules from: “The molecule is a egfr inhibitor and modulate, belonging to the cancer treatment class of molecules, and is treatment of disorder” obtained from L+M-24 test split. Note that the molecule with sufficient binding affinity to EGFR can satisfy this description. Thus, we evaluate the generated molecule using a docking tool (García-Ortegón et al., 2022). We accept the generated molecule if it yields a Vina score (Trott and Olson, 2010) below $-7.0$ when docked with an EGFR protein, which is the threshold for a potential inhibitor of the protein. We present the results in Table 12. One can see that BS generates structurally similar molecules while our method generates structurally diverse molecules.
Table 12: Visualization of the generated molecules. We generate $50$ molecules from: “The molecule is a egfr inhibitor and modulate, belonging to the cancer treatment class of molecules, and is treatment of disorder”. The blue line indicates the molecule that follows the given description, as evaluated using the external tool (docking tool; Trott and Olson, 2010). We also evaluate structurally similar molecules (Tanimoto similarity $>0.7$).
$50$ molecules generated from BS
$50$ molecules generated from ours
Appendix F Use of AI Assistants
We used AI-based writing assistants to improve the sentence structure and grammar. These tools were used only for editorial improvements. The technical content, methodology, and experimental results were entirely authored by the researchers.